twice a number is 4 more than 3 times another number. if 2 less than 7 times the second number, what are the two numbers?
Answer by mananth(16949)

twice a number is 4 more than 3 times another number. if 2 less than 7 times the second number, what are the two numbers?
one number =x
other number = y
..
2x= 3y+4
2x-3y=4...............1
7y-2=2x
7y-2x=2..............2
multiply (1)by 1
Multiply (2) by 1
2x-3y=4
-2x+7y =2

Add the two equations 0
4y=6
/4

y=1.5
Plug value of y in
2x-3y=4
2x-4.5= 4
2x=8.5
/2
x=4.25